# Hallucination Detection in LLMs with Topological Divergence on Attention Graphs

### v2

@@ 4 Experiments
We compare TOHA with seven baselines: perplexity Ren et al. (2023), max entropy Fadeeva et al. (2024), Haloscope Du et al. (2024), LLM-Check Sriramanan et al. (2024), semantic entropy Farquhar et al. (2024), EigenScore Chen et al. (2024), and SelfCheckGPT Manakul et al. (2024). Appendix E provides information on implementation details.
Table 1: ROC AUC ()($\uparrow$)( ↑ ) of hallucination detection techniques for three LLMs. The best results for each model are highlighted in bold, and the second best are underlined.
@@
SelfCheckGPT [1] 0.62 $\pm$ 0.03 0.63 $\pm$ 0.04 0.93 $\pm$ 0.02 0.82 $\pm$ 0.03 0.71 $\pm$ 0.03
Semantic entropy [2] 0.54 $\pm$ 0.06 0.57 $\pm$ 0.04 0.84 $\pm$ 0.02 0.71 $\pm$ 0.05 0.65 $\pm$ 0.05
EigenScore [3] 0.58 $\pm$ 0.02 0.55 $\pm$ 0.04 0.73 $\pm$ 0.02 0.54 $\pm$ 0.04 0.57 $\pm$ 0.03
Haloscope [4] 0.56 $\pm$ 0.05 0.54 $\pm$ 0.05 0.84 $\pm$ 0.02 0.97 $\pm$ 0.03 0.59 $\pm$ 0.05
LLM-Check [5] 0.5 $\pm$ 0.06 0.57 $\pm$ 0.03 0.62 $\pm$ 0.04 0.56 $\pm$ 0.06 0.57 $\pm$ 0.03
Perplexity [6] 0.67 $\pm$ 0.03 0.62 $\pm$ 0.04 0.77 $\pm$ 0.01 0.45 $\pm$ 0.06 0.66 $\pm$ 0.03
Max entropy [7] 0.66 $\pm$ 0.05 0.6 $\pm$ 0.05 0.73 $\pm$ 0.02 0.73 $\pm$ 0.04 0.71 $\pm$ 0.03
TOHA (ours) 0.76 $\pm$ 0.04 0.63 $\pm$ 0.04 0.89 $\pm$ 0.02 0.84 $\pm$ 0.01 0.61 $\pm$ 0.03
@@
SelfCheckGPT [1] 0.6 $\pm$ 0.02 0.62 $\pm$ 0.05 0.78 $\pm$ 0.04 0.59 $\pm$ 0.05 0.70 $\pm$ 0.06
Semantic entropy [2] 0.54 $\pm$ 0.05 0.54 $\pm$ 0.04 0.76 $\pm$ 0.03 0.63 $\pm$ 0.06 0.63 $\pm$ 0.05
EigenScore [3] 0.54 $\pm$ 0.02 0.52 $\pm$ 0.04 0.68 $\pm$ 0.03 0.53 $\pm$ 0.05 0.63 $\pm$ 0.06
Haloscope [4] 0.53 $\pm$ 0.06 0.48 $\pm$ 0.02 0.74 $\pm$ 0.04 0.58 $\pm$ 0.04 0.58 $\pm$ 0.08
LLM-Check [5] 0.46 $\pm$ 0.02 0.49 $\pm$ 0.03 0.6 $\pm$ 0.03 0.58 $\pm$ 0.05 0.58 $\pm$ 0.08
Perplexity [6] 0.66 $\pm$ 0.01 0.57 $\pm$ 0.04 0.73 $\pm$ 0.03 0.57 $\pm$ 0.09 0.58 $\pm$ 0.05
Max entropy [7] 0.66 $\pm$ 0.03 0.57 $\pm$ 0.05 0.71 $\pm$ 0.04 0.64 $\pm$ 0.02 0.56 $\pm$ 0.05
TOHA (ours) 0.67 $\pm$ 0.02 0.57 $\pm$ 0.05 0.88 $\pm$ 0.04 0.9 $\pm$ 0.04 0.63 $\pm$ 0.02
@@
SelfCheckGPT [1] 0.57 $\pm$ 0.04 0.60 $\pm$ 0.04 0.86 $\pm$ 0.04 0.78 $\pm$ 0.03 0.61 $\pm$ 0.05
Semantic entropy [2] 0.61 $\pm$ 0.03 0.51 $\pm$ 0.05 0.75 $\pm$ 0.05 0.72 $\pm$ 0.02 0.64 $\pm$ 0.04
EigenScore [3] 0.56 $\pm$ 0.04 0.49 $\pm$ 0.04 0.45 $\pm$ 0.04 0.51 $\pm$ 0.04 0.55 $\pm$ 0.02
Haloscope [4] 0.55 $\pm$ 0.05 0.50 $\pm$ 0.01 0.65 $\pm$ 0.04 0.54 $\pm$ 0.03 0.59 $\pm$ 0.01
LLM-Check [5] 0.44 $\pm$ 0.04 0.57 $\pm$ 0.04 0.54 $\pm$ 0.03 0.53 $\pm$ 0.05 0.63 $\pm$ 0.04
Perplexity [6] 0.65 $\pm$ 0.02 0.59 $\pm$ 0.02 0.58 $\pm$ 0.04 0.5 $\pm$ 0.05 0.58 $\pm$ 0.05
Max entropy [7] 0.58 $\pm$ 0.03 0.55 $\pm$ 0.06 0.70 $\pm$ 0.06 0.78 $\pm$ 0.04 0.56 $\pm$ 0.04
TOHA (ours) 0.69 $\pm$ 0.01 0.54 $\pm$ 0.04 0.93 $\pm$ 0.03 0.92 $\pm$ 0.02 0.66 $\pm$ 0.05
Table 2: ROC AUC ()($\uparrow$)( ↑ ) of hallucination detection techniques. The best results for each model are highlighted in bold, and the second best are underlined.
Method Single gen. SQuAD CoQA XSum
@@
SelfCheckGPT [1] 0.79 $\pm$ 0.05 0.76 $\pm$ 0.07 0.81 $\pm$ 0.02
Semantic entropy [2] 0.58 $\pm$ 0.04 0.83 $\pm$ 0.05 0.49 $\pm$ 0.05
EigenScore [3] 0.52 $\pm$ 0.07 0.82 $\pm$ 0.06 0.49 $\pm$ 0.05
Haloscope [4] 0.85 $\pm$ 0.02 0.56 $\pm$ 0.07 0.55 $\pm$ 0.04
LLM-Check [5] 0.48 $\pm$ 0.04 0.5 $\pm$ 0.08 0.56 $\pm$ 0.03
Perplexity [6] 0.82 $\pm$ 0.02 0.69 $\pm$ 0.04 0.61 $\pm$ 0.04
Max entropy [7] 0.5$\pm$ 0.04 0.53 $\pm$ 0.05 0.47 $\pm$ 0.01
TOHA (ours) 0.85 $\pm$ 0.02 0.73 $\pm$ 0.01 0.63 $\pm$ 0.04
@@
SelfCheckGPT [1] 0.62 $\pm$ 0.06 0.84 $\pm$ 0.04 0.74 $\pm$ 0.04
Semantic entropy [2] 0.68 $\pm$ 0.03 0.77 $\pm$ 0.06 0.65 $\pm$ 0.04
EigenScore [3] 0.67 $\pm$ 0.04 0.66 $\pm$ 0.09 0.51 $\pm$ 0.04
Haloscope [4] 0.59 $\pm$ 0.05 0.71 $\pm$ 0.05 0.58 $\pm$ 0.05
LLM-Check [5] 0.5 $\pm$ 0.05 0.54 $\pm$ 0.1 0.55 $\pm$ 0.05
Perplexity [6] 0.63 $\pm$ 0.06 0.67 $\pm$ 0.08 0.65 $\pm$ 0.05
Max entropy [7] 0.73 $\pm$ 0.05 0.74 $\pm$ 0.05 0.52 $\pm$ 0.08
TOHA (ours) 0.8 $\pm$ 0.02 0.69 $\pm$ 0.05 0.69 $\pm$ 0.02
Main results.
The results of our experiments are provided in Tables 12. We evaluate TOHA against state-of-the-art hallucination detection methods and demonstrate its competitive performance, consistently securing first or second place across most benchmark datasets. TOHA significantly outperforms uncertainty-based baselines and matches the accuracy of consistency-based approaches, achieving a notable $13.4\%$ improvement on the challenging MS MARCO dataset that includes long and detailed model responses. While SelfCheckGPT emerges as TOHA’s closest competitor, it relies on additional generations, incurring substantial computational overhead. Comparison of the best-performing baselines (Figure 6) confirms that TOHA reduces inference time by an order of magnitude compared to methods of similar quality.
Compared to Haloscope Du et al. (2024), which operates with limited annotated data and large-scale unlabeled data, TOHA not only delivers superior performance but also eliminates the need for unannotated generations for hyperparameter tuning and a separate classifier training. This makes TOHA less data-dependent and more practical for real-world applications.
Another interesting comparison is with LLM-Check Sriramanan et al. (2024), which also uses attention maps to compute hallucination scores—specifically, by averaging the log determinant of attention maps from a single pre-selected layer. However, our TOHA achieves superior performance, demonstrating that not all attention heads contribute equally to hallucination detection. By employing a topology-based head selection strategy, we significantly enhance detection quality.
Generalizability to different data distributions.
From a deployment perspective, hallucination detection methods must remain robust to shifts in input data distribution, given the inherent diversity of real-world user queries. To evaluate TOHA’s robustness in this regard, we conducted transfer experiments on Mistral-7B (see Figure 4a). The results highlight TOHA’s strong transferability: for the XSum and CNN/DM datasets, performance changes in transfer settings fall within the standard deviation. For the remaining datasets, TOHA maintains competitive performance compared to baseline methods (Table 1), demonstrating its adaptability to diverse data distributions.
@@
As previously mentioned, TOHA requires only a small set of samples to identify ”hallucination-aware" attention heads. To assess its sensitivity to probe set size, we conducted an ablation study (Figure 4). The results demonstrate TOHA’s robustness to limited annotated data: even with just $50$ samples, performance does not drop significantly and mostly remains stable as the probe set size increases.
Figure 4: (a) Generalizability between the datasets, model: Mistral-7B-Instruct. The vertical axis corresponds to the origin of the probe set, the horizontal axis  to the test dataset. (b)-(c): Detection quality dependence on the size of a probe set, models: Mistral-7B (left), LLama-2-7B (right).
What do hallucination patterns look like?
As detailed in Section 3, the topological divergences we employ characterize the MSF connecting the vertices $R$ of the response to the vertices $P$ of the prompt. For hallucination-aware heads, we analyzed MSF patterns distinguishing hallucinated and grounded samples. A key finding is that hallucinated samples frequently exhibit strong attention to the <s> token, whereas grounded samples tend to attend to <s> less (Figure 5).
To verify the significance of <s> in hallucination detection, we conducted an ablation experiment: after removing <s> from the texts, we recomputed the TOHA hallucination scores for the selected hallucination-aware heads. The results (Table 3) show a significant performance drop, confirming that attention to <s> is a critical indicator of hallucination. This finding aligns with prior work demonstrating the influential role of <s> in LLM mechanisms Barbero et al. (2025). However, using the average attention to <s> alone as a hallucination score proves insufficient (see Table 9 in Appendix D). In contrast, our proposed score, which incorporates the intricate structure of attention maps, demonstrates far greater discriminative power for this task.
Table 3: TOHA performance with and without <s> token. TOP-1 results are highlighted with bold font.
Dataset with <s> w/o <s>
Mistral-7B
MS MARCO 0.67 $\pm$ 0.02 0.56 $\pm$ 0.02
CoQA 0.88 $\pm$ 0.02 0.32 $\pm$ 0.03
LLaMA-2-7B
MS MARCO 0.76 $\pm$ 0.04 0.66 $\pm$ 0.03
CoQA 0.89 $\pm$ 0.02 0.56 $\pm$ 0.04
Figure 5: Attention to <s>: a) hallucinated sample and b) grounded one. Green color denotes edges and nodes that correspond to grounded tokens of a generation, yellow color — hallucinated ones. Model: Mistral-7B.
Figure 6: Comparison of methods’ inference time in seconds. The measurements were obtained for $16$ random samples from MS MARCO, model: Mistral-7B. For the SelfCheckGPT, $20$ additional answers were generated for each sample.
5 Related works
Hallucination detection methods.
The problem of hallucinations in LLMs has attracted significant attention recently Zhang et al. (2023); Huang et al. (2023); Wang et al. (2024). Consistency-based methods  Manakul et al. (2024); Chen et al. (2024); Kuhn et al. (2023); Qiu and Miikkulainen (2024); Nikitin et al. (2024) that use the diversity of multiple LLM responses as a hallucination score offer robust detection but impose significant computational overhead. Surface-level techniques like perplexity and logit entropy Fadeeva et al. (2024); Malinin and Gales (2021) analyze model confidence directly from output distributions — efficient but limited in detection capability as they neglect the model’s rich internal representations. Hidden states-based classifiers Azaria and Mitchell (2023); Sky et al. (2024); Zhou et al. (2025) require extensive annotated datasets, which are scarce in the public domain Zhang et al. (2023). This issue was partially addressed by the Haloscope Du et al. (2024), which leverages unlabeled data “in the wild” with minimum annotated data needed for hyperparameter selection. Attention map-based methods represent a promising yet underdeveloped direction. Current techniques either rely on large labeled data, e.g., Lookback Lens Chuang et al. (2024), or exploit only simple attention graph properties, such as self-loop weights in LLM-Check Sriramanan et al. (2024). This leaves a critical research gap: training-free methods that fully leverage the rich structural information encoded in attention relationships remain underexplored.
Topological Data Analysis (TDA) in NLP.
Topological Data Analysis (TDA) is a mathematical framework for extracting multi-scale structural patterns from data using principles from topology and computational geometry Chazal and Michel (2017); Hensel et al. (2021). Recent years have seen growing interest in applying TDA to natural language processing (NLP) tasks to study textual structural properties Uchendu and Le (2024). For example,  Tulchinskii et al. (2024) leveraged persistent homology to estimate the intrinsic dimensionality of CLS embeddings for detecting machine-generated text. Other work has demonstrated the utility of topological features derived from transformer attention matrices — treated as weighted graphs — for diverse NLP applications. These include uncertainty quantification Kostenok et al. (2023) and grammatical acceptability classification Proskurina et al. (2023b), where topological features extracted from the attention graphs were used as input to train auxiliary classifiers.

### v3

@@ 4.1 Experiment Setting
We compare TOHA with a comprehensive set of eight baselines: uncertainty-based perplexity Ren et al. (2023) and max entropy Fadeeva et al. (2024); inner states-based ReDEEP Sun et al. (2025), HaloScope Du et al. (2024), and LLM-Check Sriramanan et al. (2024); consistency-based semantic entropy Farquhar et al. (2024), EigenScore Chen et al. (2024), and SelfCheckGPT Manakul et al. (2024).
Implementation details.
The reported results are averaged over 5 runs with different data splits, using test sets comprising 25% of the data and a fixed validation set size of 100, following the methodology of HaloScope Du et al. (2024). To ensure a fair comparison, we consider two implementations of the latter: a standard setting with $20$ generations (see Tables 5-6), and an efficiency-comparable one with minimal number of generations — $1$ for SelfCheckGPT, $5$ for semantic entropy and EigenScore (Tables 1-2). Appendix F provides additional information on implementation details.
Table 1: ROC AUC $(\uparrow)$ of hallucination detection techniques for three LLMs. The best results for each model are highlighted in bold, and the second best are underlined.
@@
SelfCheckGPT [1] 0.63 $\pm$ 0.04 0.51 $\pm$ 0.04 0.86 $\pm$ 0.02 0.71 $\pm$ 0.04 0.66 $\pm$ 0.04
Semantic entropy [2] 0.54 $\pm$ 0.03 0.51 $\pm$ 0.04 0.83 $\pm$ 0.02 0.70 $\pm$ 0.03 0.56 $\pm$ 0.03
EigenScore [3] 0.54 $\pm$ 0.04 0.50 $\pm$ 0.06 0.74 $\pm$ 0.02 0.71 $\pm$ 0.04 0.58 $\pm$ 0.04
HaloScope [4] 0.57 $\pm$ 0.08 0.51 $\pm$ 0.10 0.62 $\pm$ 0.08 0.92 $\pm$ 0.07 0.62 $\pm$ 0.02
LLM-Check [5] 0.49 $\pm$ 0.03 0.49 $\pm$ 0.03 0.60 $\pm$ 0.01 0.50 $\pm$ 0.03 0.58 $\pm$ 0.04
Perplexity [6] 0.45 $\pm$ 0.01 0.54 $\pm$ 0.02 0.54 $\pm$ 0.03 0.81 $\pm$ 0.05 0.54 $\pm$ 0.06
Max entropy [7] 0.68 $\pm$ 0.04 0.60 $\pm$ 0.07 0.73 $\pm$ 0.00 0.75 $\pm$ 0.05 0.71 $\pm$ 0.02
ReDEEP [8] 0.54 $\pm$ 0.02 0.47 $\pm$ 0.06 0.59 $\pm$ 0.03 0.45 $\pm$ 0.05 0.63 $\pm$ 0.04
TOHA (ours) 0.76 $\pm$ 0.04 0.60 $\pm$ 0.09 0.89 $\pm$ 0.01 0.96 $\pm$ 0.01 0.66 $\pm$ 0.05
@@
SelfCheckGPT [1] 0.59 $\pm$ 0.03 0.60 $\pm$ 0.03 0.66 $\pm$ 0.03 0.57 $\pm$ 0.03 0.64 $\pm$ 0.05
Semantic entropy [2] 0.53 $\pm$ 0.03 0.51 $\pm$ 0.03 0.76 $\pm$ 0.01 0.73 $\pm$ 0.03 0.61 $\pm$ 0.04
EigenScore [3] 0.55 $\pm$ 0.03 0.53 $\pm$ 0.04 0.61 $\pm$ 0.03 0.75 $\pm$ 0.02 0.63 $\pm$ 0.02
HaloScope [4] 0.51 $\pm$ 0.05 0.48 $\pm$ 0.05 0.61 $\pm$ 0.04 0.67 $\pm$ 0.04 0.57 $\pm$ 0.07
LLM-Check [5] 0.44 $\pm$ 0.02 0.49 $\pm$ 0.06 0.60 $\pm$ 0.03 0.49 $\pm$ 0.01 0.61 $\pm$ 0.01
Perplexity [6] 0.54 $\pm$ 0.04 0.44 $\pm$ 0.03 0.74 $\pm$ 0.02 0.46 $\pm$ 0.03 0.56 $\pm$ 0.09
Max entropy [7] 0.65 $\pm$ 0.04 0.59 $\pm$ 0.06 0.65 $\pm$ 0.03 0.73 $\pm$ 0.04 0.56 $\pm$ 0.03
ReDEEP [8] 0.54 $\pm$ 0.04 0.52 $\pm$ 0.04 0.72 $\pm$ 0.04 0.42 $\pm$ 0.08 0.54 $\pm$ 0.06
TOHA (ours) 0.65 $\pm$ 0.02 0.56 $\pm$ 0.02 0.90 $\pm$ 0.01 0.87 $\pm$ 0.04 0.68 $\pm$ 0.05
@@
SelfCheckGPT [1] 0.58 $\pm$ 0.04 0.58 $\pm$ 0.05 0.77 $\pm$ 0.02 0.64 $\pm$ 0.03 0.60 $\pm$ 0.04
Semantic entropy [2] 0.57 $\pm$ 0.04 0.54 $\pm$ 0.03 0.76 $\pm$ 0.04 0.65 $\pm$ 0.03 0.60 $\pm$ 0.03
EigenScore [3] 0.56 $\pm$ 0.04 0.47 $\pm$ 0.04 0.57 $\pm$ 0.03 0.57 $\pm$ 0.02 0.52 $\pm$ 0.06
HaloScope [4] 0.54 $\pm$ 0.09 0.51 $\pm$ 0.04 0.57 $\pm$ 0.03 0.55 $\pm$ 0.02 0.55 $\pm$ 0.07
LLM-Check [5] 0.49 $\pm$ 0.06 0.56 $\pm$ 0.05 0.57 $\pm$ 0.02 0.57 $\pm$ 0.07 0.57 $\pm$ 0.07
Perplexity [6] 0.54 $\pm$ 0.04 0.46 $\pm$ 0.07 0.62 $\pm$ 0.03 0.45 $\pm$ 0.02 0.49 $\pm$ 0.05
Max entropy [7] 0.62 $\pm$ 0.03 0.53 $\pm$ 0.06 0.66 $\pm$ 0.03 0.78 $\pm$ 0.02 0.59 $\pm$ 0.04
ReDEEP [8] 0.62 $\pm$ 0.06 0.48 $\pm$ 0.05 0.73 $\pm$ 0.02 0.48 $\pm$ 0.07 0.58 $\pm$ 0.08
TOHA (ours) 0.67 $\pm$ 0.04 0.56 $\pm$ 0.05 0.92 $\pm$ 0.02 0.88 $\pm$ 0.05 0.66 $\pm$ 0.03
4.2 Results
Main results.
The results of our experiments are provided in Tables 12. We evaluate TOHA against state-of-the-art hallucination detection methods and demonstrate its competitive performance TOHA significantly outperforms uncertainty-based baselines and matches the quality of consistency-based approaches, achieving notable improvements of $11.7\%$ on the challenging long-form QA MS MARCO dataset for Mistral-7B and $21.6\%$ on the conversational QA dataset CoQA for LLama-2-7B.
To evaluate TOHA’s robustness to data distribution change, we conducted transfer experiments on Mistral-7B, see Figure 3(c) for the results. TOHA possesses strong transferability: for the XSum and CNN/DM datasets, performance in transfer settings fall within the standard deviation of the method (see Table 9), while remaining competitive on the other datasets (Table 1).
Evaluation on a multi-hop dataset.
To validate TOHA in a more realistic setting, we consider an additional experiment using the HotpotQA Yang et al. (2018) dataset. It consists of questions that require knowledge from multiple supporting documents — much like real-world queries, which also rarely rely on a single source of information. The results are provided in Table 3. TOHA demonstrates superior performance to baselines, confirming its effectiveness “in the wild”.
Efficiency.
Figure 4(a) shows that TOHA is approximately seven times faster than SelfCheckGPT with a single additional generation. Given that SelfCheckGPT typically requires 10-20 generations, this makes TOHA over 70 times faster in practice. Furthermore, TOHA’s runtime is close to the lightweight entropy baseline; with a more efficient low-level implementation, it has the potential to approach the cost of a single forward pass while achieving higher accuracy than other inexpensive alternatives.
Table 2: ROC AUC $(\uparrow)$ of hallucination detection techniques. The best results for each model are highlighted in bold, and the second best are underlined.
Method Single gen. CoQA SQuAD XSum
@@ 4.2 Results
SelfCheckGPT [1] 0.91 $\pm$ 0.01 0.65 $\pm$ 0.05 0.68 $\pm$ 0.05
Semantic entropy [2] 0.78 $\pm$ 0.02 0.55 $\pm$ 0.07 0.47 $\pm$ 0.04
EigenScore [3] 0.80 $\pm$ 0.02 0.45 $\pm$ 0.05 0.50 $\pm$ 0.03
HaloScope [4] 0.67 $\pm$ 0.08 0.84 $\pm$ 0.07 0.55 $\pm$ 0.06
LLM-Check [5] 0.54 $\pm$ 0.06 0.49 $\pm$ 0.05 0.52 $\pm$ 0.04
Perplexity [6] 0.51 $\pm$ 0.05 0.68 $\pm$ 0.02 0.65 $\pm$ 0.03
Max entropy [7] 0.82 $\pm$ 0.03 0.60 $\pm$ 0.02 0.65 $\pm$ 0.03
ReDEEP [8] 0.58 $\pm$ 0.08 0.39 $\pm$ 0.04 0.62 $\pm$ 0.06
TOHA (ours) 0.84 $\pm$ 0.01 0.87 $\pm$ 0.03 0.65 $\pm$ 0.05
@@
SelfCheckGPT [1] 0.69 $\pm$ 0.01 0.74 $\pm$ 0.02 0.69 $\pm$ 0.02
Semantic entropy [2] 0.76 $\pm$ 0.04 0.57 $\pm$ 0.04 0.64 $\pm$ 0.03
EigenScore [3] 0.78 $\pm$ 0.03 0.55 $\pm$ 0.04 0.55 $\pm$ 0.04
HaloScope [4] 0.66 $\pm$ 0.13 0.75 $\pm$ 0.04 0.57 $\pm$ 0.07
LLM-Check [5] 0.53 $\pm$ 0.07 0.54 $\pm$ 0.06 0.54 $\pm$ 0.02
Perplexity [6] 0.39 $\pm$ 0.03 0.65 $\pm$ 0.03 0.53 $\pm$ 0.05
Max entropy [7] 0.85 $\pm$ 0.02 0.47 $\pm$ 0.05 0.60 $\pm$ 0.06
ReDEEP [8] 0.37 $\pm$ 0.10 0.56 $\pm$ 0.03 0.67 $\pm$ 0.03
TOHA (ours) 0.79 $\pm$ 0.05 0.77 $\pm$ 0.02 0.69 $\pm$ 0.03
Figure 3: (a)-(b): Detection quality dependence on the size of a probe set, models: Mistral-7B (left), LLama-2-7B (right). (c) Generalizability between the datasets, model: Mistral-7B. The vertical axis corresponds to the origin of the probe set, the horizontal axis to the test dataset.
Method Single gen. Mistral-7B Llama-2-13B
SelfCheckGPT 0.70 $\pm$ 0.06 0.63 $\pm$ 0.04
Semantic entropy 0.70 $\pm$ 0.05 0.70 $\pm$ 0.04
EigenScore 0.68 $\pm$ 0.04 0.67 $\pm$ 0.04
HaloScope 0.60 $\pm$ 0.06 0.50 $\pm$ 0.01
LLM-Check 0.48 $\pm$ 0.03 0.56 $\pm$ 0.04
Perplexity 0.55 $\pm$ 0.06 0.49 $\pm$ 0.04
Max entropy 0.62 $\pm$ 0.04 0.69 $\pm$ 0.05
ReDEEP 0.49 $\pm$ 0.04 0.62 $\pm$ 0.05
TOHA (ours) 0.71 $\pm$ 0.08 0.80 $\pm$ 0.03
Table 3: ROC-AUC $(\uparrow)$ values on the HotpotQA dataset. Best results are highlighted in bold, and the second best are underlined.
4.3 Analysis of Hallucination-Aware Attention Heads
Attention patterns and copying behaviour.
For hallucination-aware heads, we analyzed minimal spanning forest (MSF) patterns distinguishing hallucinated and grounded samples. A key finding is that for these heads, hallucinated samples frequently exhibit strong attention to the first token, whereas grounded samples tend to attend to the first token less. An example is provided in Figure 5. Since this pattern is a known behavior of token induction, or copying heads — which default to the first token when unable to find previous occurrences of the current token in the context Elhage et al. (2021) — we decided to explore the relationship between these special heads and hallucination-aware ones.
Figure 4: a): Inference time comparison (seconds) for various methods evaluated on 16 MS MARCO samples using Mistral-7B. SelfCheckGPT measurement includes one additional generated answer per sample. b)-c): ROC-AUC performance of TOHA across different numbers of selected attention heads ($N_{max}$) on Mistral-7B and Llama-2-7B.
Figure 5: Attention to the first token (<s> in this example) for (a) a hallucinated generation and (b) a grounded one. Green highlights edges and nodes corresponding to grounded tokens, while yellow indicates hallucinated tokens. Model: Mistral-7B.
Figure 6: Copying ranks of hallucination-aware attention heads (lower ranks indicate stronger copying behaviour). The row labels $(X,Y)$ correspond to $X$-th layer and $Y$-th head.
To investigate this link, we ranked all attention heads in Llama-2-7B and Mistral-7B based on their copying scores, following the method of Feucht et al. (2025). For the subset of heads frequently identified by TOHA (appearing in $\geq 20\%$ of runs across all datasets), their copying ranks were recorded. The results, shown in Figure 6, reveal that hallucination-aware heads are often also among the model’s top-25 copiers, a finding that aligns with prior work Sun et al. (2025).
4.4 Ablation Studies
Why zero out the distances between prompt tokens?
We hypothesize that the connections within the prompt contribute little to hallucination detection; therefore, our method is designed to disregard them. To validate this architectural choice, we considered an ablation study, where $\operatorname{MTop-Div}_{G}(R,P)$ was replaced by the MST length of the complete graph in Algorithm 1. The results showing the effectiveness of the proposed approach are in Table 4.
Table 4: ROC-AUC $(\uparrow)$ of Algorithm 1 with MST length of the complete graph vs $\operatorname{MTop-Div}_{G}(R,P)$. Best results are highlighted in bold.
Dataset MST length $\operatorname{MTop-Div}_{G}(R,P)$
Mistral-7B
CoQA $0.75\pm 0.03$ $\mathbf{0.90\pm 0.01}$
MS MARCO $0.38\pm 0.03$ $\mathbf{0.65\pm 0.02}$
LLama-2-7B
CoQA $0.60\pm 0.03$ $\mathbf{0.89\pm 0.01}$
MS MARCO $0.37\pm 0.02$ $\mathbf{0.76\pm 0.04}$
What about other attention map-based features?
To demonstrate that our topology-based approach offers an advantage over conventional attention map features, we compared them to $\operatorname{MTop-Div}$ in the supervised setting. Table 8 from Appendix B shows that a classifier trained on $\operatorname{MTop-Div}_{G}(R,P)$ achieves the best performance, confirming that our score captures unique patterns in an attention map that standard approaches miss.
@@ 4.4 Ablation Studies
To evaluate TOHA’s sensitivity to probe set size, we conducted a sensitivity study. The results from Figure 3 confirm TOHA’s robustness: even with just $50$ samples, the performance drop remains minor with stable results as the probe set size increases.
How many attention heads do we need?
To evaluate the sensitivity of our method to the hyperparameter $N_{max}$, we performed an ablation study for values from $1$ to $10$. As shown in Figure 4, TOHA achieves strong detection performance even when $N_{\max}=1$, which underscores the effectiveness of our topological approach.